Tuition of $1669 will be due when the spring term begins in 4 months. What amount should a student deposit today, at 3.35 %, to have enough to pay the tuition?


1
Expert's answer
2020-09-08T15:23:31-0400

Tuition of $1,669 will be due when the spring term begins in 4 months.


The amount that a student should pay today at 3.35% will be ;


Present Value (PV) ="\frac{Future Value (FV) } {(1+r)^{t}}"


PV = "\frac{FV} {(1 + r)^{4}}"


="\frac{1669}{(1+\frac{0.0335}{12})^{4}}"


=$1650.49